uppose milk and honey are the only products produced in Ghana. Use the data below from the
economy of Ghana to answer the questions that follow:
Year Price of Milk Quantity of Milk Price of Honey Quantity of Honey
2011 Ghc1 100 Ghc2 50
2012 Ghc1 200 Ghc2 100
2013 Ghc2 200 Ghc4 100
Base year: 2011
Required:
a. Compute for each year the
i. nominal GDP, (3marks)
ii. real GDP, (3marks)
iii. the GDP deflator. (3marks)
b. Compute for 2012 and 2013 from the respective preceding years the percentage change in
i. nominal GDP, (2marks)
ii. real GDP, (2marks)
iii. the GDP deflator (2marks)
1
Expert's answer
2021-08-02T11:07:14-0400

a.

i. Nominal GDP is:

NGDP(2011) = 1×100 + 2×50 = 200,

NGDP(2012) = 1×200 + 2×100 = 400,

NGDP(2013) = 2×200 + 4×100 = 800.

ii. Real GDP is:

RGDP(2011) = 1×100 + 2×50 = 200,

RGDP(2012) = 1×200 + 2×100 = 400,

RGDP(2013) = 1×200 + 2×100 = 400.

iii. The GDP deflator is:

2011: 200/200×100 = 100,

2012: 400/400×100 = 100,

2013: 800/400×100 = 200.

b. In 2012 and 2013 the percentage change in

i. nominal GDP is:

2012: (400/200 - 1)×100% = 100%,

2013: (800/400 - 1)×100% = 100%.

ii. real GDP is:

2012: (400/200 - 1)×100% = 100%,

2013: (400/400 - 1)×100% = 0%.

iii. the GDP deflator is:

2012: (100/100 - 1)×100% = 0%,

2013: (200/100 - 1)×100% = 100%.